A franchise of automotive technicians has a contract arrangement for a rewards program with a car manufacturing company calleneed the correct amounts
Answer #1

Total selling price = 5 +0.45 = 5.45

Transaction price allocated to Automotive services = 23,000,000 * 5/5.45 = 21,100,917.4

Transaction price allocated to Automotive services = 23,000,000 * 0.45/5.45 = 1,899,082.6

Any doubt please comment, If it has helped you, please give a thumbs up.